# MathFusion: Enhancing Mathematic Problem-solving of LLM through Instruction Fusion

### v1

@@ Appendix D Effect of Teacher Model
In MathFusion, we use GPT-4o-mini (OpenAI et al., 2023) as the teacher model to generate the solutions for the fused problems. To validate the performance improvement of MathFusion is not merely due to the stronger teacher model, we conduct an ablation study, where we use GPT-4o-mini to rewrite the solutions from the original training set. Then we fine-tune the Llama3-8B model on the original training set and the rewritten solutions. The results are shown in Table 6. We can see that the performance of the model fine-tuned on the rewritten solutions is better than the Standard setting, especially on the MATH and GSM8K datasets. However, the average improvement is only 1.3 points. Meanwhile, each fusion strategy of MathFusion still outperforms the rewritten solution by a large margin, indicating that the performance improvement of MathFusion mainly comes from the fusion of problems rather than the stronger teacher model.
Appendix E Significant Test
@@
Appendix F More Cases
More cases, including the original problems and , the fused problem , are shown below. Specifically, we show three reasonable cases in Case F, Case F, and Case F, and three unreasonable cases in Case F, Case F, and Case F.

### v2

@@ Appendix D Effect of Teacher Model
In MathFusion, we use GPT-4o-mini (OpenAI et al., 2023) as the teacher model to generate the solutions for the fused problems. To validate the performance improvement of MathFusion is not merely due to the stronger teacher model, we conduct two ablation studies: (1) use GPT-4o-mini to rewrite the solutions from the original training set; and (2) follow DART-Math (Tong et al., 2024) to use DeepSeekMath-7B-RL (Shao et al., 2024) to generate solutions for the fused problems. The results are shown in Table 6. We can see that the performance of the model fine-tuned on the rewritten solutions is better than the Standard setting, especially on the MATH and GSM8K datasets. However, the average improvement is only 2.9 points. Meanwhile, each fusion strategy of MathFusion still outperforms the rewritten solution by a large margin. Additionally, though DeepSeekMath-7B-RL underperforms GPT-4o-mini in distillation quality (34.6 v.s. 36.1 on average), it still outperforms DART-Math (34.6 v.s. 32.9 on average). These results indicate that the performance improvement of MathFusion mainly comes from the new problem generated by three fusion strategies rather than the stronger teacher model.
Appendix E Additional Baseline
A most recent work, Mosaic-IT (Li et al., 2024c), shares similar idea with our MathFusion. Mosaic-IT is a model-free data augmentation technique that operates by concatenating existing instruction-following datasets and subsequently training LLMs using these augmented data instances along with meta-instructions. We conduct a comparison with the “Primary Strategy” proposed in Mosaic-IT, where the question pairs (same as MathFusion) and corresponding solutions from the original GSM8K and MATH datasets are concatenated into a single sample for SFT, resulting in 15K data. To mitigate overfitting to the pattern of answering multiple questions jointly, we also conduct an additional experiment which combine Mosaic-IT and the original GSM8K and MATH training sets during training, resulting in 30K data in total. The results are shown in Table 6. We observe that the Mosaic-IT leads to inferior performance, even worse than the Standard setting (i.e., using only the original GSM8K and MATH training data). We suspect this may be due to the lack of logical integration between problems when simply concatenated—unlike MathFusion, which explicitly introduces semantic or reasoning connections (e.g., sequential dependency or comparative logic) through its fusion strategies. This highlights the advantage of model-driven, structure-aware fusion over model-free concatenation.
Appendix F Significant Test
@@
Appendix G More Cases
More cases, including the original problems and , the fused problem , are shown below. Specifically, we show three reasonable cases in Case G, Case G, and Case G, and three unreasonable cases in Case G, Case G, and Case G.